Ok. I need to know how to set this word problem up into an equation. "Roger has three times as many apples as Sam. and Sam has twice as many apples as Christi. How many apples does each have if Roger has 12 more than both Sam and Christi together?"
Answer by scott8148(6628)

three unknowns means three equations

"Roger has three times as many apples as Sam" ___ R = 3S

"Sam has twice as many apples as Christi" ___ S = 2C

"Roger has 12 more than both Sam and Christi together" ___ R = S + C + 12

substituting ___ 3S = S + C + 12

substituting ___ 6C = 2C + C + 12 ___ 3C = 12 ___ C = 4

substituting ___ S = 2(4) = 8

substituting ___ R = 3(8) = 24